OpenStudy (anonymous):

What is 5.8 × 106 mg/s in grams per minute? A. 9.7 × 109 g/min B. 9.7 × 107 g/min C. 3.48 × 105 g/min D. 3.48 × 103 g/min

OpenStudy (anonymous):

exponents

OpenStudy (anonymous):

9.7 x10^9

OpenStudy (anonymous):

can anyone help me

OpenStudy (jdoe0001):

\(\bf \cfrac{5.8\times 10^6\cancel{ mg }}{\cancel{ sec }}\cdot \cfrac{g}{1000\cancel{ mg }}\cdot \cfrac{60\cancel{ sec}}{min}\)

OpenStudy (imstuck):

You can set it up like this: